1. The pendulum has the largest kinetic energy at the lowest point of its swing. TRUE or FALSE

2.      At what launch angle does a projectile go the farthest before it hits the ground

90

0

30

45

60

30 and 60

3. When you are loading your furniture into a moving truck, you will do less work if you use a shorter (and steeper) ramp, TRUE or FALSE
Answer #1

1. TRUE

Highest point of swing No kinetic energy Maximum gravitational potential energy Maximum kinetic energy Minimum gravitational

2. 45 degrees

3.FALSE
Answer #2

1.) false

2) 45

3.) false
Answer #3

1. TRUE

Highest point of swing No kinetic energy Maximum gravitational potential energy Maximum kinetic energy Minimum gravitational

2. 45 degrees

The basic motion equation x=Voxt Range of Trajector)y can be used to find the range By symmetry, the total time of flight is

3.FALSE. The work against gravity is same and is equal to mgh. where h is the height of the truck and m is the mass of furniture
Answer #4

u=0 Fr= max FT = min Us = max KE = 0 Fr = max KE-0 FT min θ max θ=0 u=max U,=min Fr=0 KE-max FT max 0 KE max

45 degrees goes the farthest. This is due to the nature of right triangles. Additionally, from the equation for the range :

u sin 2θ

We can see that the Range will be maximum when the value of \sin 2 \theta is the highest (i.e, when it is equal to 1). Clearly, 2 \theta has to be 90 degrees. That is to say, \theta is 45 degrees.

A) TRUE

B) 45